Skyler is 6 years older than Destiny. In 4 years the sum of their ages will be 44. How old is Skyler now?
Answer by ikleyn(52803)

.
Skyler is 6 years older than Destiny. In 4 years the sum of their ages will be 44. How old is Skyler now?
~~~~~~~~~~~~~~~~~~


Let Skyler age be x.

Then the Destiny age is (x-6) years.



Next, write equations for their ages in 4 years, as you read the problem

    (x+4) + ((x-6)+4) = 44.


Simplify and find x

    x + 4 + x - 6 + 4 = 44

    2x +2 = 44

    2x    = 44 - 2

    2x = 42

     x = 42/2 = 21.


ANSWER.  Skyler is 21 years old.


CHECK.  (a) Destiny is 21-6 = 15 years old.

        (b)  In 4 years the sum of their ages will be  (21+4) + (15+4) = 25 + 19 = 44.    ! Correct !

Solved.